OpenStudy (meilendurcer):

Tell which number is prime: 15, 24, 41, 62. A. 15 B. 41 C. 62 D. 24

hartnn (hartnn):

15=3*5 62=31*2 24=6*4 does this help ?

OpenStudy (anonymous):

15 can be divided by 3, 24 by 2, 62 by 2, so 41 is the prime